OpenStudy (anonymous):

NEED HELP I WILL GIVE A MEDAL FOR THE RIGHT ANSWER !!! Find the coordinates for the center for this circle. (x + 3)² + (y - 5)² = 25 (3,- 5) (-3, 5) (-5, 3)

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\]the equation of a circle with center at (h, k) and radius r. h = the number you put in for x to make the inside = 0 for yours, x+3 is inside so what value of x makes it 0? that's h. same method/ideology goes for k.

OpenStudy (anonymous):

So number 1 ?

OpenStudy (anonymous):

what's number 1?

OpenStudy (anonymous):

(3,- 5)

OpenStudy (anonymous):

well... put in 3 for x. does 3 + 3 = 0?

OpenStudy (anonymous):

well no , so my second guess would be (-3, 5) ?

OpenStudy (anonymous):

why would you guess? look, set x+3 = 0 and slove for x. that will be the x coordinate of your center. do the same with y, i.e., set y-5=0 and solve for y. that will be the y coordinate of your center. do it like that and you'll never have to guess!